MsBrains (ms-brains):

PLEASE HELP ME WITH 5 Q'S! I WILL MEDAL!

MsBrains (ms-brains):

1. I think it is B or C not sure though

OpenStudy (fortytherapper):

It is one of those. Which do you think?

MsBrains (ms-brains):

I don't know...I'm leaning more in C.. But honestly I suck at these kinda of problems.

OpenStudy (fortytherapper):

The equation is y-y1 = m(x-x1) What would happen if you did Y - (-1)? What would you do to those two minus signs?

MsBrains (ms-brains):

Make them in to + signs?

OpenStudy (fortytherapper):

Exactly! Same with the x's X - (-3)

MsBrains (ms-brains):

Okay! So it is B then. cx Yay. Can you help me with 4 more?

OpenStudy (fortytherapper):

I'll see what I can do. I'm on my phone without paper waiting to go to work But let's see them

MsBrains (ms-brains):

Not sure about this one... Thanks

OpenStudy (fortytherapper):

Lets see if I can do this without paper So, does Ax + By = C look familiar?

MsBrains (ms-brains):

Not really..

OpenStudy (fortytherapper):

Idk yet, ima work it out in my head real quick, so I may write random things here to keep track

MsBrains (ms-brains):

Okay. cx

OpenStudy (fortytherapper):

So this was the same graph and was Y+1 = 6/7(X+3) So first you have to distribute the 6/7 into the X and 3

OpenStudy (fortytherapper):

Y+1 = 6/7x + 18/7

OpenStudy (fortytherapper):

Then lets move the y 1 = 6/7x - Y + 18/7

OpenStudy (fortytherapper):

1 is 7/7, so subtract 18/7 -11/7 = 6/7x - Y I think And then

OpenStudy (fortytherapper):

Yeah, that's the furthest I can get without paper and ima have to head in, sorry and good luck!

MsBrains (ms-brains):

No probs. Thanks dude. :)